# MedS3: Towards Medical Small Language Models with Self-Evolved Slow Thinking

### v1

@@ 5.2 Inference-Time Scaling
In this section, we discuss whether MedS3 can benefit from more inference tokens. We sample $n=2,4,8,12,16$ samples for a given problem, and conduct PRM-Vote-Sum to select the most plausible response. We select five clinical datasets (Healthfact, DDXPlus, SEER, BioMRC, Pubhealth) to illustrate the scaling law, and plot the results in Fig. 3. We observe that the first 4 samples provide significant performance improvements, showing a satisfactory tradeoff between the tokens count and the performance gain. Although more generations bring less significant performance gains, the increasing trend never slows down except for the Pubhealth dataset, illustrating unbounded scaling potentials. For the sake of large-scale evaluation, we finalize the search budget to 16 across all experiments.
6 Related Works
@@
7 Future Work
@@
8 Conclusion
In this paper, we present MedS3, a self-evolved slow-thinking medical language model built for universal clinical usage. We collect a small seed dataset, with only 7,000 instances but covering 16 difference realistic medical tasks, and use Monte-Carlo Tree Search to construct policy data and PRM data. We propose a new decoding method, which enables the resulting policy model to collaborate with the fine-tuned PRM model, to produce credible long-chain responses. Experiment results demonstrate that our model achieves state-of-the-art performance on eleven downstream medical benchmarks, especially in realistic clinical ones, even surpassing GPT-4o-mini with a large margin.
Appendix A Prompt Template
@@
Appendix B Dataset Details
In this section, we elucidate the seed dataset and the evaluation sets:
@@ Appendix B Dataset Details
    PubmedQA (Jin et al., 2019) is a specialized benchmark for biomedical question answering, consisting of question-answer pairs derived from PubMed abstracts. It focuses on yes/no/maybe questions that require reasoning over biomedical literature. We use the human-labeled question set and split the training set and test set, with both 500 problems for evolution and evaluation, respectively.
@@
Appendix C Hyperparameters
C.1 Data Synthesis
@@
C.2 Self-Training of Policy and PRM
@@
We use PRMTrainer of trl to train the PRM model. We use the same hyper-parameters as those used in the policy fine-tuning but conduct a simple data filtering method. For an instance where all its solution trajectories are all correct or incorrect, we only randomly maintain one of them, as the excessively simple or hard instance will disturb the training process. For other instances, we guarantee that the correct trajectories and incorrect trajectories have the same total count for a given data point, which performs a simple 0-1 class balance. We train 1 epoch for filtered 76,792 samples.
C.3 Evaluation
For evaluation, the temperature is set to $1.0$ and top_p is set to $0.9$. The max generation tokens are set to 8,192. We use exact match (EM) to measure the performance. Specifically, we extract the contents following “The answer is” template, and perform appropriate post-process to derive a final prediction. For multiple-choice problems, we directly choose the first character of prediction phrases and measure whether the ground truth is equal to the prediction. For close-ended generation tasks, we remove quotes and turn the prediction and the ground truth into lowercase phrases. After that, we check whether the ground truth phrases exist in the prediction phrases.
Appendix D Case Study

### v2

@@ 5.2 Inference-Time Scaling
In this section, we discuss whether MedS3 can benefit from more inference tokens. We sample $n=2,4,8,12,16$ candidates for a given problem, and conduct P-VS to select the most plausible response. We select five clinical datasets (Healthfact, DDXPlus, SEER, BioMRC, Pubhealth) to illustrate the scaling law, and plot the results in Fig. 4. We observe that the first 4 samples provide significant performance improvements, showing a satisfactory tradeoff between the tokens consumption and the performance improvement. Although more generations bring fewer significant performance gains, the increasing trend never slows down except for the Pubhealth dataset, illustrating nearly unbounded scaling potentials.
Figure 5: Three widely adopted methods to empower models with medical reasoning abilities. MCTS+PRM is the best among the three, making it the core of MedS3.
5.3 Comparison of Slow-Thinking Styles
In this section, we compare MCTS plus PRM which is what MedS3 leverages, with distillation from strong reasoning models, which is what O1-journey-part3 (Huang et al., 2025) does and pure reinforcement learning (RL), which is what DeepSeek-R1 (Guo et al., 2025) adopts. These methodologies are widely used for empowering small language models with strong long-chain reasoning abilities. We use the same dataset in §3 to implement RL, and use the officially released distillation dataset provided by Huang et al. (2025) to SFT the base model. The results presented in Fig. 5 demonstrate that in exam-level medical QA datasets where the base model already excels at, distillation from large proprietary reasoning models is much more data-efficient than the other two methods, albeit sacrificing generalization in clinical tasks. In contrast, with both considerable performance leap and generalization, RL is second to MCTS+PRM. We hypothesize that the medical diagnosis step is easier to determine than math reasoning steps, resulting in a more accurate PRM.
6 Conclusion
In this paper, we present MedS3, a self-evolved slow-thinking medical language model built for universal clinical usage. We collect a seed dataset, with only 7,465 instances but covering 16 different realistic medical tasks, and use Monte-Carlo Tree Search to construct policy data and PRM data. We propose a new decoding method, which enables the resulting policy model to collaborate with the fine-tuned PRM model, to produce credible long-chain responses. Experiment results demonstrate that our model achieves superior performance on eleven downstream medical benchmarks, especially in realistic clinical ones, surpassing open-sourced models with a large margin.
Limitations
MedS3 achieves superior performance over eleven benchmarks by conducting MCTS in seed datasets to collect both policy and PRM training data and a newly proposed decoding strategy: PRM-guided Vote-Sum. However, it can be further improved via these strategies: (1) conduct an iterative self-evolution pipeline to enhance both the policy and PRM; (2) cooperate with reinforcement learning to empower the policy with the “aha-moment” (Guo et al., 2025) ability; (3) introduce more training samples to avoid data imbalance.
Ethics Considerations
In developing clinical slow-thinking model MedS3, it is crucial to address ethical consideration that arise when utilizing AI in healthcare environments. Below are the key ethical considerations that have been taken into account:
Perfomance vs. Potential Risks
While MedS3 demonstrates significant enhancements in clinical reasoning and task performance, it is important to acknowledge the inherent limitations of AI models. These models can generate misleading information of “hallucinations”, which could pose risks in clinical settings. Therefore, MedS3 is not intended to replace medical professionals or provide definitive clinical decisions but rather to assist healthcare providers under appropriate supervision.
Data Ethics and Privacy Compliance
All patient data has been anonymized, and informed consent was obtained for its use, ensuring full compliance with privacy policies and obtaining explicit permission for all data usage. Additionally, data usage has been approved by relevant ethics committees, ensuring compliance with ethical standards and privacy protection requirements.
Appendix A Related Works
@@
Appendix B Further Experiments
In this section, we present more experiments to validate the effectiveness of MedS3.
B.1 Soft Train vs. Hard Train
We here compare the PRM training techniques, where the step label is set to a hard two-class $0/1$ label (Zhang et al., 2025) or a soft float label introduced in MedS3. We remain the other settings unchanged and set the label to $1$ if the step’s rollout value is greater than 0, otherwise 0 when training the hard PRM. The results in Table 4 indicate that soft-label is comprehensively superior to hard-label. Although Zhang et al. (2025) indicate that hard-label works after prompting an LLM to filter bad intermediate steps, for the sake of privacy in clinical uasge and fairness we do not conduct filtering.
B.2 Reliability of PRM
Although our process reward model (PRM) is trained using rollout-value—a rule-based supervision signal—we empirically demonstrate that the PRM exhibits strong capability in identifying erroneous reasoning steps. To evaluate this, we compare two metrics: the task error rate (defined as instances where model predictions deviate from ground truth answers) and the PRM error rate (occurring when the PRM assigns a higher score to an incorrect candidate than to the ground truth-aligned prediction). As illustrated in Table 5, the majority of tasks exhibit a PRM error rate of zero, with the exception of the Healthfact dataset. This suggests that our fine-tuned PRM effectively differentiates between valid and invalid clinical reasoning steps. Furthermore, the observed discrepancy between task and PRM error rates implies that most errors arise not from the PRM’s assessment but from the policy model’s failure to generate candidates aligned with the ground truth. This finding underscores the need to refine the policy model in future iterations to address this limitation.
Method MedQA MedMCQA Healthfact DDXPlus SEER Average
Hard 64.81 62.18 77.61 64.35 54.94 64.778
Soft 71.88 65.20 79.97 65.20 58.36 68.122
Table 4: Comparison between soft-label which is adopted in MedS3 and hard-label.
Error Type
MedQA
MedMCQA HealthFact PubmedQA BioASQ Med MMLU BioMRC DDXPlus PubHealth DrugDose SEER
Task 28.91 34.60 20.03 40.40 23.79 20.50 22.88 34.60 26.97 4.35 41.87
PRM 0.00 0.00 16.16 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Table 5: Comparison of error rates (lower is better) in task-level and PRM-level, where PRM works with no errors in most of testbeds.
Appendix C Future Work
@@
Appendix D Prompt Template
@@
Appendix E Dataset Details
In this section, we elucidate the seed dataset and the evaluation sets. We divide the used 16 training datasets into the following five dimensions:
  1. 1.
    Long Context QA: This dimension enables MedS3 to capture useful information from the given context and response with long-chain reasoning. This dimension covers BioMRC (Pappas et al., 2020), HeadQA Topic Classification (Vilares and Gómez-Rodríguez, 2019; Wu et al., 2024b), and DDX-Plus (Tchango et al., 2022).
  2. 2.
    Knowledge-Intensive QA: This dimension teaches MedS3 to use long-chain reasoning to answer knowledge-intensive problems, which covers MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), PubMedQA (Jin et al., 2019) and PubHealth (Kotonya and Toni, 2020).
  3. 3.
    Bio-Medical QA: This part leverages general data in bio-medicine domains to enhance the generality of MedS3, which includes SciQ (Welbl et al., 2017), Evidence Inference (DeYoung et al., 2020) and Head QA (Vilares and Gómez-Rodríguez, 2019).
  4. 4.
    Medical Natural Language Inference: This dimension prompts MedS3 to discriminate biomedical research concepts and corresponding descriptions, which contains Healthfact Classification (Kotonya and Toni, 2020), Medical Question Pair (MQP; McCreery et al. (2020)), and catalonia-independence-corpus (CIC; Zotova et al. (2020)).
  5. 5.
    Diagnosis QA: This dimension is related to real-world clinical scenarios, including disease diagnosis and classification and drug related questions. We choose Covid-19 Classification (Lab, 2020), Drug-Dose Extraction and Adverse Drug Event Classification (Huynh et al., 2016; Wu et al., 2024b).
The descriptions of each training and evaluation datasets are presented below:
@@ Appendix E Dataset Details
    PubmedQA (Jin et al., 2019) is a specialized benchmark for biomedical question answering, consisting of question-answer pairs derived from PubMed abstracts. It focuses on yes/no/maybe questions that require reasoning over biomedical literature. We use the human-labeled question set and split the training set and test set, with both 500 problems for evolution and evaluation, respectively. Note that we do not include relevant contexts before questions, challenging models’ internal knowledge comprehension.
@@
Appendix F Hyperparameters
F.1 Data Synthesis
@@
F.2 Self-Training of Policy and PRM
@@
We use PRMTrainer of trl to train the PRM model. We use the same hyper-parameters as those used in the policy fine-tuning but conduct a simple data filtering method. For an instance where all its solution trajectories are all correct or incorrect, we only randomly maintain one of them, as the excessively simple or hard instance will disturb the training process. For other instances, we guarantee that the correct trajectories and incorrect trajectories have the same total count for a given data point, which performs a simple 0-1 class balance. We train 1 epoch for filtered 76,792 samples with LoRA, where the rank $r$ and $\alpha$ are set to 64 and 128, respectively.
F.3 Evaluation
For evaluation, the temperature is set to $1.0$ and top_p is set to $0.9$. The max generation tokens are set to 8,192. Among the three presented decoding mechanisms, CoT (Wei et al., 2022) directly prompts models to generate a long reasoning chain and outputs the answer with “The answer is {answer}” for the convenience of answer extraction. Self-Consistency (Wang et al., 2023) generates $N=16$ samples for a given problem, and we select the one whose answer appears most times among the $N$ outputs. For PRM-guided Vote-Sum, we generate $N=12$ samples for datasets where SC has already achieved a high score (MedMCQA, Med-MMLU and BioASQ) and generate $N=16$ samples for other datasets. This is reasonable since not all generated samples are of high quality due to the sampling parameters top_p. We manually filter some low-probability samples to prevent them from disturbing the Vote-Sum computation. We use exact match (EM) to measure the performance. Specifically, we extract the contents following the last “The answer is” template to match the self-reflection thinking style, and perform appropriate post-process to derive a final prediction. For multiple-choice problems, we directly choose the first character of prediction phrases and measure whether the ground truth is equal to the prediction. For close-ended generation tasks, we remove quotes and turn the prediction and the ground truth into lowercase phrases. After that, we check whether the ground truth phrases exist in the prediction phrases.
F.4 Training Details of Distillation and RL
In this section, we elucidate the implementation details of distillation and RL.
Distillation
[…]
RL
We follow Guo et al. (2025) to use Group Relative Policy Optimization (GRPO; Shao et al. (2024)) to conduct RL training. We set the number of generations to 10 and the learning rate to $61e-$ 6. We adopt ZeRO-3 (Rajbhandari et al., 2020) to save memory and conduct full fine-tuning in one 8xA100 machine. The batch size is set to 4 per GPU. For the adopted prompt, we use the same prompt illustrated in DeepSeek-R1-zero, and use <think></think><answer></answer> to learn the slow-thinking output style. We use accuracy reward and format reward, and the reward setting is presented below in Table 6.
Table 6: Reward configurations for training RL-style medical models. Null means that there is no contents between <answer> and </answer>.
Conditions Accuracy Format
Correct 1 1
Incorrect -1 -1
Null 0 Null
Appendix G AI Assistant Usage
We mainly use AI assistants to polish our expressions.
Appendix H Case Study